# 1Neural Networks and Deep Learning Lab, MIPT, Dolgoprudny, Russia

### v1

@@ 2 Associative Recurrent Memory Transformer
$\overline{v}_{i}=\frac{A_{i-1}^{l}\phi(k_{i})}{(z_{i-1})^{T}\phi(k_{i})};\quad\gamma=1-\frac{(z_{i-1})^{T}\phi(k_{i})}{\|\phi(k_{i})\|^{2}};$ (2)
$A_{i}^{l}=A_{i-1}^{l}+\beta_{i}(v_{i}-\overline{v}_{i})\otimes\phi(k_{i});\quad z^{l}_{i}=z^{l}_{i-1}+\gamma\phi(k_{i}).$ (3)
@@
Note that without $\gamma$ ($\gamma=1$) this approach suffers from catastrophic forgetting on some tasks. The reason is that while we erase the information $\overline{v}_{i}$ from the $A$-matrix, the corresponding keys remain in the normalization vector $z_{i}$. As shown in our experiments (LABEL:fig:rewrite_ablation), this problem becomes significant when performing hundreds of erase-insert operations with associative memory. To overcome this, we propose to take into account the previous keys in $z_{i}$ when updating it (details are in Section F.1).
@@ F.1 Gamma-correction
$\gamma=1-\frac{(z_{i-1})^{T}\phi(k_{i})}{\|\phi(k_{i})\|^{2}};$
$z_{i}=z_{i-1}+\gamma\phi(k_{i}).$
The inner product $(z_{i-1})^{T}\phi(k_{i})$ is divided by the square of $\|\phi(k_{i})\|$ because the gamma will be further multiplied by $\phi(k_{i})$.

### v2

@@ 2 Associative Recurrent Memory Transformer
$\overline{v}_{i}=\frac{A_{s-1}^{l}\phi(k_{i})}{(z_{s-1})^{T}\phi(k_{i})};\quad\gamma_{i}=1-\frac{(z_{s-1})^{T}\phi(k_{i})}{\|\phi(k_{i})\|^{2}};$ (2)
$A_{s}^{l}=A_{s-1}^{l}+\sum_{i}\beta_{i}(v_{i}-\overline{v}_{i})\otimes\phi(k_{i});\quad z^{l}_{s}=z^{l}_{s-1}+\sum_{i}\gamma_{i}\phi(k_{i}).$ (3)
@@
Note that without $\gamma_{i}$ ($\gamma_{i}=1$) this approach suffers from catastrophic forgetting on some tasks. The reason is that while we erase the information $\overline{v}_{i}$ from the $A$-matrix, the corresponding keys remain in the normalization vector $z_{s}$. As shown in our experiments (LABEL:fig:rewrite_ablation), this problem becomes significant when performing hundreds of erase-insert operations with associative memory. To overcome this, we propose to take into account the previous keys in $z_{s}$ when updating it (details are in Section F.1).
@@ F.1 Gamma-correction
$\gamma_{i}=1-\frac{(z_{s-1})^{T}\phi(k_{i})}{\|\phi(k_{i})\|^{2}};$
$z_{s}=z_{s-1}+\sum_{i}\gamma_{i}\phi(k_{i}).$
The inner product $(z_{s-1})^{T}\phi(k_{i})$ is divided by the square of $\|\phi(k_{i})\|$ because the gamma will be further multiplied by $\phi(k_{i})$.